# GenAI Arena: An Open Evaluation Platform for Generative Models

### v3

@@ Abstract
Generative AI has made remarkable strides to revolutionize fields such as image and video generation. These advancements are driven by innovative algorithms, architecture, and data. However, the rapid proliferation of generative models has highlighted a critical gap: the absence of trustworthy evaluation metrics. Current automatic assessments such as FID, CLIP, FVD, etc often fail to capture the nuanced quality and user satisfaction associated with generative outputs. This paper proposes an open platform GenAI-Arena to evaluate different image and video generative models, where users can actively participate in evaluating these models. By leveraging collective user feedback and votes, GenAI-Arena aims to provide a more democratic and accurate measure of model performance. It covers three arenas for text-to-image generation, text-to-video generation, and image editing respectively. Currently, we cover a total of 27 open-source generative models. GenAI-Arena has been operating for four months, amassing over 6000 votes from the community. We describe our platform, analyze the data, and explain the statistical methods for ranking the models. To further promote the research in building model-based evaluation metrics, we release a cleaned version of our preference data for the three tasks, namely GenAI-Bench. We prompt the existing multi-modal models like Gemini, GPT-4o to mimic human voting. We compute the accuracy by comparing the model voting with the human voting to understand their juding abilities. Our results show existing multimodal models are still lagging in assessing the generated visual content, even the best model GPT-4o only achieves an average accuracy of $49.19\%$ across the three generative tasks. Open-source MLLMs perform even worse due to the lack of instruction-following and reasoning ability in the complex vision scenarios.
@@ 1 Introduction
Since February 11th, 2024, we have collected over 6000 votes for three multimodal generative tasks. We constructed leaderboards for each task with these votes, identifying the state-of-the-art models as PlayGround V2.5, MagicBrush, and T2VTurbo, respectively (until June 4th, 2024). Detailed analyses based on the votes are presented. For example, our plotted winning fraction heatmaps reveal that while the Elo rating system is generally effective, it can be biased by imbalances between "easy games" and "hard games". We also performed several case studies for qualitative analysis, demonstrating that users can provide preference votes from multiple evaluation aspects, which help distinguish subtle differences between the outputs and upload high-quality votes for Elo rating computation.
@@ 3.1 Design
GenAI-Arena is designed to offer an intuitive and comprehensive evaluation platform for generative models, facilitating user interaction and participation. The platform is structured around three primary tasks: text-to-image generation, image edition, and text-to-video generation. Each task is supported by a set of features that include an anonymous side-by-side voting system, a battle playground, a direct generation tab, and a leaderboard as shown in Figure 2 . These features are designed to cater to both casual users and researchers, ensuring a democratic and accurate assessment of model performance.
@@
To ensure a fair comparison between different models, we ported the highly dispersed codebase from the existing works and then standardized them into a unified format. During inference, we fixed the hyper-parameters and the prompt format to prevent per-instance prompt or hyper-parameter tuning, which makes the inference of different models fair and reproducible. Following ImagenHub [31], we build the new library of VideoGenHub, which aims to standardize the inference procedure for different text-to-video and image-to-video models. We find the best hyper-parameters of these models to ensure their highest performance.
@@ 3.2 Model Integration
In Table 2, we list all the included text-to-image generation models. For example, SDXL, SDXL-Turbo, and SDXL-Lightning are all derived based on SDXL [59], while SDXL-Turbo [65] and SDXL-Lightning [43] adopt different distillation method. We also include diffusion transformer models [56] like PixArt-$\alpha$ and PixArt-$\sigma$. Playground V2 and Playground V2.5 are based on SDXL architecture, but trained by Playground.ai from scratch with an internal dataset.
@@
SDXL [59] 3.5B Latent Diffusion Model 1K×$\times$1K 50
SDXL-Turbo [65] 3.5B Latent Diffusion Model + Distillation 1K×$\times$1K 1
SDXL-Lightning [43] 3.5B Latent Diffusion Model + Distillation 1K×$\times$1K 4
PixArt-$\alpha$ [9] 0.6B Diffusion Transformer 1K×$\times$1K 50
PixArt-$\sigma$ [10] 0.6B Diffusion Transformer + Weak-to-Strong 4K×$\times$4K 50
StableCascade [58] 1.5B + 3.6B Würstchen Architecture 1K×$\times$1K 20+10
Playground V2 [39] 3.5B Latent Diffusion Model 1K×$\times$1K 50
Playground V2.5 [38] 3.5B Latent Diffusion Model 1K×$\times$1K 50
@@
AnimateDiff [16] SD-1.5 2s 8 WebVid10M 512 x 512 25
AnimateDiff-Turbo [16] SD-1.5 2s 8 WebVid10M 512 x 512 4
ModelScope [68] SD-1.5 2s 8 WebVid10M 256 x 256 50
LaVie [70] SD-1.5 2s 8 Vimeo25M 320 x 512 50
StableVideoDiffusion [4] SD-2.1 2.5s 10 LVD-500M 576 x 1024 20
VideoCrafter2 [8] SD-2.1 2s 16 WebVid10M 320 x 512 50
T2V-Turbo [40] VideoCrafter2 2s 8 WebVid10M 320 x 512 4
OpenSora [51] Pixart-$\alpha$ 2s 16 WebVid10M 320 x 512 50
@@ 3.3 Elo Rating System
where $S(i,j)$ is the actual match outcome, $S(i,j)=1$ for a win $S(i,j)=0.5$ for a tie, and $S(i,j)=0$ for a loss, and $E(i,j)=P(Y{i,j}=1)$.
@@
$\mathcal{L}(\mathbf{R})=\sum_{i,j\in N,i\neq j}\left(W_{ij}Y_{ij}\log P(Y_{ij}=1)\right)$ (3)
@@ 3.4 GenAI-Museum
To achieve this, we serve GenAI-Museum as a pre-computed data pool comprising various inputs from existing datasets or user collection, along with each model’s output. Based on this, a "Random Sample" button shown in  Figure 2 is additionally implemented to facilitate the random generation of prompts and the immediate retrieval of corresponding images or videos. This functionality operates by sending requests to our deployed GenAI-Museum every time "Random Sample" button is hit, receiving input and two random model’s pre-computed outputs. In this way, we save the computation time on the GPU, enable users to do instant comparisons and votes on the UI, and balance the votes for each unique input so we gradually collect votes for a full combination of all models. The input prompts were sampled from ImagenHub [31] and VBench [24].
@@
(Last updated on June 4th, 2024)
(a) Text-to-Image
@@
PlayGround V2.5 1150 +21/-22
Playground V2 1101 +14/-20
StableCascade 1057 +20/-24
SDXL-Lightning 1053 +22/-22
PixArt-$\alpha$ 1052 +15/-19
PixArt-$\sigma$ 1050 +26/-23
SDXL 1001 +15/-14
SDXL-Turbo 935 +18/-16
OpenJourney 853 +13/-17
LCM 817 +20/-20
@@
MagicBrush $1$ +28/-32
InfEdit 1079 +27/-33
CosXLEdit 1066 +31/-30
InstructPix2Pix 1033 +32/-26
PNP 998 +37/-36
Prompt2prompt 988 +25/-25
CycleDiffusion 939 +23/-26
SDEdit 929 +25/-21
Pix2PixZero 857 +20/-24
@@
T2V-Turbo 1113 +53/-46
StableVDiffusion 1105 +45/-37
VideoCrafter2 1077 +18/-18
AnimateDiff 1075 +23/-26
LaVie 997 +24/-26
OpenSora 916 +19/-23
ModelScope 866 +19/-22
AnimateDiff-Turbo 851 +18/-20
@@
4 Benchmarks and Results Discussion
4.1 Arena Leaderboard
We report our leaderboard at the time of paper writing in Table 5. For image generation, we collected 4443 votes in total. The currently top-ranked models are Playground V2.5 and Playground V2. Both of the models are released by Playground.ai, which follows the same architecture as SDXL but is trained with a private dataset. In contrast, SDXL only ranks in the seventh position, lagging significantly behind. Such finding highlights the importance of the training dataset. Following the Playground models is StableCascade, which utilizes a highly efficient cascade architecture to lower the training cost. According to Würstchen [58], StableCascade only requires a 10% training cost of SD-2.1, yet it can beat SDXL significantly on our leaderboard. This highlights the importance of the diffusion architecture to achieve strong performance. For image editing, a total of 1083 votes have been collected. MagicBrush, InFEdit, CosXLEdit, and InstructPix2Pix ranked higher as they can perform localized editing on images. PNP preserves the structure with feature injections, thus limiting the edit variety. The older methods such as Prompt-to-Prompt, CycleDiffusion, SDEdit, and Pix2PixZero, frequently result in completely different images during editing despite the high-quality images, which explains the lower ranking of these models. For text-to-video, there is a total of 1568 votes. T2VTurbo leads with the highest Elo score, suggesting it is the most effective model. Close behind, StableVideoDiffusion ranks second. Following VideoCrafter2 and AnimateDiff have very close elo scores, showing nearly equivalent capabilities. LaVie, OpenSora, ModelScope, and AnimateDiff-Turbo follow with decreasing scores, indicating progressively lower performance.
@@ 4.2 Discussion and Insights
Specific cells in the heatmap reveal notable findings. For instance, although PlayGround 2.5 achieves the state-of-the-art (SOTA) Elo rating in the Text-to-Image task, its winning fraction over PixArt-$\sigma$ is only $0.48$, which is below 50%. Similarly, the Text-to-Video SoTA model, T2V-Turbo, has a lower winning fraction against StableVideoDiffusion. The higher Elo rating of T2V-Turbo might be due to our Arena collecting more votes from "easy games" with low-ranked models and fewer from "harder games" with high-ranked models. For example, the number of battles between T2V-Turbo and AnimateDiff-Turbo ($30$) is way more than T2V-Turbo with other models (around $10$) in  Figure 4. These anomalies highlight potential drawbacks of the Elo rating system: (1) a reliable and robust Elo rating requires a large amount of voting data, and (2) the estimated Elo rating may be biased by the imbalance between "easy games" and "harder games," as they carry similar weight in the estimation.
@@ 6 Conclusion
In this paper, we introduced GenAI-Arena, an open platform designed to rank generative models across text-to-image, image editing, and text-to-video tasks based on user preference. unlike other platforms, GenAI-Arena is driven by community voting to ensure transparency and sustainable operation. We employed the side-by-side human voting method to evaluate the models and collected over 6000 votes starting from February 11th, 2024. We compiled an Elo leaderboard with the votings and found that PlayGround V2.5, MagicBrush, and T2V-Turbo are the current state-of-the-art models in the three tasks (until June 4th, 2024). Analysis based on the collected votes shows that while the Elo rating is generally functional, but can biased by the imbalance of the "easy games" and "hard games". Several case studies demonstrate the high quality of our collected votes. What’s more, we also released the human preference voting as GenAI-Bench. We prompt the existing MLLMs to evaluate the generated images and videos on GenAI-Bench and compute the accuracy with human voting. The experiment showed that the open-source MLLMs achieve very low performance, even the best model GPT-4o can only achieve $49.19\%$ accuracy. This is mostly because their lack of instruction-following and reasoning ability in complex vision scenarios. In the future, we will continue collecting human votes to update the leaderboard, helping the community to keep track of the research progress. We also plan to develop a more robust MLLM to better approximate human ratings in GenAI-Bench.

### v4

@@ Abstract
Generative AI has made remarkable strides to revolutionize fields such as image and video generation. These advancements are driven by innovative algorithms, architecture, and data. However, the rapid proliferation of generative models has highlighted a critical gap: the absence of trustworthy evaluation metrics. Current automatic assessments such as FID, CLIP, FVD, etc often fail to capture the nuanced quality and user satisfaction associated with generative outputs. This paper proposes an open platform GenAI-Arena to evaluate different image and video generative models, where users can actively participate in evaluating these models. By leveraging collective user feedback and votes, GenAI-Arena aims to provide a more democratic and accurate measure of model performance. It covers three tasks of text-to-image generation, text-to-video generation, and image editing respectively. Currently, we cover a total of 35 open-source generative models. GenAI-Arena has been operating for seven months, amassing over 9000 votes from the community. We describe our platform, analyze the data, and explain the statistical methods for ranking the models. To further promote the research in building model-based evaluation metrics, we release a cleaned version of our preference data for the three tasks, namely GenAI-Bench. We prompt the existing multi-modal models like Gemini, and GPT-4o to mimic human voting. We compute the accuracy by comparing the model voting with the human voting to understand their judging abilities. Our results show existing multimodal models are still lagging in assessing the generated visual content, even the best model GPT-4o only achieves an average accuracy of $49.19\%$ across the three generative tasks. Open-source MLLMs perform even worse due to the lack of instruction-following and reasoning ability in complex vision scenarios.
@@ 1 Introduction
Since February 11th, 2024, we have collected over 9000 votes for three multimodal generative tasks. We constructed leaderboards for each task with these votes, identifying the state-of-the-art models as PlayGround V2.5, MagicBrush, and StableVideoDiffusion, respectively (until Oct 24th, 2024). Detailed analyses based on the votes are presented. For example, our plotted winning fraction heatmaps reveal that while the Elo rating system is generally effective, it can be biased by imbalances between "easy games" and "hard games". We also performed several case studies for qualitative analysis, demonstrating that users can provide preference votes from multiple evaluation aspects, which help distinguish subtle differences between the outputs and upload high-quality votes for Elo rating computation.
@@ 3.1 Design
GenAI-Arena is designed to offer an intuitive and comprehensive evaluation platform for generative models, facilitating user interaction and participation. The platform is structured around three primary tasks: text-to-image generation, image edition, and text-to-video generation. Each task is supported by a set of features that include an anonymous and a non-anonyumous battle playground, a direct generation tab, and a leaderboard as shown in Figure 2 . These features are designed to cater to both casual users and researchers, ensuring a democratic and accurate assessment of model performance.
@@
To ensure a fair comparison between different models, we ported the highly dispersed codebase from the existing works and then standardized them into a unified format. During inference, we fixed the hyper-parameters and the prompt format to prevent per-instance prompt or hyper-parameter tuning, which makes the inference of different models fair and reproducible. Following ImagenHub [32], we build the new library of VideoGenHub (details in  subsection A.5), which aims to standardize the inference procedure for different text-to-video and image-to-video models. We find the best hyper-parameters of these models to ensure their highest performance.
@@ 3.2 Model Integration
In Table 2, we list all the included text-to-image generation models. For example, SDXL, SDXL-Turbo, and SDXL-Lightning are all derived based on SDXL [63], while SDXL-Turbo [69] and SDXL-Lightning [47] adopt different distillation method. We also include diffusion transformer models [60] like PixArt-$\alpha$ and PixArt-$\sigma$. Playground V2 and Playground V2.5 are based on SDXL architecture, but trained by Playground.ai from scratch with an internal dataset. We have also included the latest released HunyuanDiT [45], FLUX.1-dev [35], FLUX.1-schnell [35].
@@
SDXL [63] 3.5B Latent Diffusion Model (LDM) 1K×$\times$1K 50
SDXL-Turbo [69] 3.5B LDM + Distillation 1K×$\times$1K 1
SDXL-Lightning [47] 3.5B LDM + Distillation 1K×$\times$1K 4
PixArt-$\alpha$ [9] 0.6B Diffusion Transformer (DiT) 1K×$\times$1K 50
PixArt-$\sigma$ [10] 0.6B DiT + Weak-to-Strong 4K×$\times$4K 50
StableCascade [62] 1.5B + 3.6B Würstchen 1K×$\times$1K 20+10
Playground V2 [42] 3.5B LDM 1K×$\times$1K 50
Playground V2.5 [41] 3.5B LDM 1K×$\times$1K 50
FLUX.1-dev [35] 12B Guidance-distilled DiT + Flow Matching 1K×$\times$1K 20
FLUX.1-schnell [35] 12B Timestep-distilled DiT + Flow Matching 1K×$\times$1K 4
Kolors [33] 2.6B LDM + ChatGLM3 1K×$\times$1K 50
HunyuanDiT [45] 1.5B DiT + multilingual text encoder 1K×$\times$1K 50
Stable Diffusion 3 [70] 8B Multimodal DiT 1K×$\times$1K 50
AuraFlow [14] 6.8B Flow-based Model 1K×$\times$1K 50
@@
AnimateDiff [17] SD-1.5 2s 8 WebVid10M 512$\times$512 25
AnimateDiff-Turbo [17] SD-1.5 2s 8 WebVid10M 512$\times$512 4
ModelScope [73] SD-1.5 2s 8 WebVid10M 256$\times$256 50
LaVie [75] SD-1.5 2s 8 Vimeo25M 320$\times$512 50
StableVideoDiffusion [4] SD-2.1 2.5s 10 LVD-500M 576$\times$1024 20
VideoCrafter2 [8] SD-2.1 2s 16 WebVid10M 320$\times$512 50
T2V-Turbo [43] VideoCrafter2 2s 8 WebVid10M 320$\times$512 4
OpenSora [55] Pixart-$\alpha$ 2s 16 WebVid10M 320$\times$512 50
OpenSora v1.2 [55] Pixart-$\alpha$ 2s 16 WebVid10M 320$\times$512 50
CogVideoX-2B [80] DiT 2s 8 35M videos + 2B images 480$\times$720 50
@@ 3.3 Elo Rating System
where $S(i,j)$ is the actual match outcome, $S(i,j)=1$ for a win $S(i,j)=0.5$ for a tie, and $S(i,j)=0$ for a loss, and $E(i,j)=P(Y_{ij}=1)$. K is
@@
$\mathcal{L}(\mathbf{R})=\sum_{i,j\in N,i\neq j}\left(W_{ij}\log P(Y_{ij}=1)\right)=\sum_{i,j\in N,i\neq j}\left(W_{ij}\log\frac{1}{1+10^{(R_{j}-R_{i})/\alpha}}\right)$ (3)
@@
Selection of battle pair
With a limited number of games, choosing which two players to match up is a crucial issue. The simplest approach, which we currently use, is to randomly select two players. However, this can introduce bias, with some models getting significantly more matches than others. A vote-aware selection system that increases the probability of selecting less-played models and lowers it for more-played ones is needed, and we plan to explore this in future Arena improvements.
@@ 3.4 GenAI-Museum
To achieve this, we serve GenAI-Museum as a pre-computed data pool comprising various inputs from existing datasets or user collection, along with each model’s output. Based on this, a "Random Sample" button shown in  Figure 2 is additionally implemented to facilitate the random generation of prompts and the immediate retrieval of corresponding images or videos. This functionality operates by sending requests to our deployed GenAI-Museum every time "Random Sample" button is hit, receiving input and two random model’s pre-computed outputs. In this way, we save the computation time on the GPU, enable users to do instant comparisons and votes on the UI, and balance the votes for each unique input so we gradually collect votes for a full combination of all models. The input prompts were sampled from ImagenHub [32] and VBench [25]. To prevent the bias in the prompt distribution, we also periodically update the input prompts with the lastest collected real-world human votes. We make sure every prompt is filtered via NSFW detector before adding them.
4 Benchmarks and Results Discussion
4.1 Arena Leaderboard
We report our leaderboard at the time of paper publishing in Table 5. For image generation, we collected 6300 votes in total. The currently top-1 model is Playground V2.5, released by Playground.ai, which follows the same architecture as SDXL but is trained with a private dataset. In contrast, SDXL only ranks in the thirteenth position, lagging significantly behind. Such finding highlights the importance of the training dataset. StableCascade is ranked in the sixth place in the leaderboard, which utilizes a highly efficient cascade architecture to lower the training cost. According to Würstchen [62], StableCascade only requires a 10% training cost of SD-2.1, yet it can beat SDXL significantly on our leaderboard. This highlights the importance of the diffusion architecture to achieve strong performance. For image editing, a total of 1154 votes have been collected. MagicBrush, InFEdit, CosXLEdit, and InstructPix2Pix ranked higher as they can perform localized editing on images. PNP preserves the structure with feature injections, thus limiting the edit variety. The older methods such as Prompt-to-Prompt, CycleDiffusion, SDEdit, and Pix2PixZero, frequently result in completely different images during editing despite the high-quality images, which explains the lower ranking of these models. For text-to-video, there is a total of 2024 votes. StableVideoDiffusion leads with the highest Elo score, suggesting it is the most effective model. Close behind, CogVideoX-2B ranks second. The following VideoCrafter2 and AnimateDiff have very close elo scores, showing nearly equivalent capabilities. LaVie, OpenSora, ModelScope, and AnimateDiff-Turbo follow with decreasing scores, indicating progressively lower performance.
@@ 4.1 Arena Leaderboard
(Last updated on Oct 24th, 2024)
(a) Text-to-Image (Top-10)
@@
PlayGround V2.5 1122 +19/-20
FLUX.1-dev 1114 +45/-42
FLUX.1-schnell 1085 +43/-46
Playground V2 1072 +18/-22
Kolors 1069 +32/-39
StableCascade 1057 +17/-21
HunyuanDiT 1030 +25/-27
PixArt-$\alpha$ 1020 +17/-19
SDXL-Lightning 1020 +17/-15
PixArt-$\sigma$ 1019 +22/-20
@@
MagicBrush 1108 +32/-28
InfEdit 1075 +26/-32
CosXLEdit 1066 +31/-29
InstructPix2Pix 1038 +32/-24
PNP 998 +35/-34
Prompt2prompt 988 +26/-23
CycleDiffusion 943 +26/-26
SDEdit 924 +24/-23
Pix2PixZero 858 +25/-30
@@
StableVideoDiffusion 1148 +31/-28
CogVideoX-2B 1106 +71/-71
T2V-Turbo 1085 +36/-32
VideoCrafter2 1068 +20/-22
AnimateDiff 1068 +25/-21
LaVie 996 +25/-21
OpenSora 912 +23/-23
OpenSora v1.2 894 +54/-71
ModelScope 862 +25/-22
AnimateDiff-Turbo 861 +22/-20
@@
(a) Text-to-Image
(b) Image Editing
(c) Text-to-Video
@@ 4.2 Discussion and Insights
Specific cells in the heatmap reveal notable findings. For instance, although PlayGround 2.5 achieves the state-of-the-art (SOTA) Elo rating in the Text-to-Image task, its winning fraction over PixArt-$\sigma$ is only $0.58$, which is below 60%. The higher Elo rating of T2V-Turbo might be due to our Arena collecting more votes from "easy games" with low-ranked models and fewer from "harder games" with high-ranked models. For example, the number of battles between PlayGround V2.5 and SDXL-Turbo ($93$) is way more than PlayGround V2.5 with other models (around $50$) in  Figure 5.
These anomalies highlight potential drawbacks of the Elo rating system: (1) a reliable and robust Elo rating requires a large amount of voting data, and (2) the estimated Elo rating may be biased by the imbalance between "easy games" and "harder games," as they carry similar weight in the estimation.
As shown in  Figure 5, we observe that the average win rates of the top-ranked models are all quite similar, none exceeding 80%. This indicates that there is no dominant, highly powerful model in text-to-image, image editing, or text-to-video generation at this time. The community is still awaiting a "ChatGPT moment"—the release of a breakthrough model with transformative capabilities.
Quality assessment of collected human votes
Since our arena users come from different backgrounds and have different preferences, we conduct an expert review on a small set of sampled human vote to ensure there are no severe quality issues of our collected votes. We let different authors review 50 items for each set. A total of 350 items from our GenAI-Bench are evaluated. During the annotations, we skipped those bad items due to NSFW or technical issues, and we finally collected 303 valid evaluations. For each vote, 3 available labels for provided for annotating:
  • Clearly Reasonable Vote: This vote will be clearly agreed by most of the people.
  • Vague Vote: The current vote makes sense. But it’s also reasonable if other vote is selected.
  • Wrong Vote: This vote will be clearly disagreed by most of the people.
Table 6: Expert Review for 350 sampled human votes
(a) Distribution of Valid Votes
# Valid votes # NSFW # Tech issue Total
303 17 30 350
86.57% 4.86% 8.57% 100%
(b) Distribution of quality labels
# Clearly Reasonable Vote # Vague Vote # Wrong Vote Total
231 51 21 303
76.24% 16.83% 6.93% 100%
We report the distribution of valid votes in  Table 6, and find that 86.57% of the votes are valid without NSFW issues. Among these valid votes, about 76.24% of the votes are clearly reasonable votes and 93.07% of the votes are either clearly reasonable or vaguely reasonable, as shown in  Table 6. We believe this shows the reliability of our preference data.
@@ 6 Conclusion
In this paper, we introduced GenAI-Arena, an open platform designed to rank generative models across text-to-image, image editing, and text-to-video tasks based on user preference. unlike other platforms, GenAI-Arena is driven by community voting to ensure transparency and sustainable operation. We employed the side-by-side human voting method to evaluate the models and collected over 9000 votes starting from February 11th, 2024. We compiled an Elo leaderboard with the votings and found that PlayGround V2.5, MagicBrush, and StableVideoDiffusion are the current state-of-the-art models in the three tasks (until Oct 24th, 2024). Analysis based on the collected votes shows that while the Elo rating is generally functional, but can biased by the imbalance of the "easy games" and "hard games". Our expert review of 350 sampled human votes confirmed that 93.07% of the votes can be viewed as either clearly reasonable or vaguely reasonable, demonstrating the high quality of our collected votes What’s more, we also released the human preference voting as GenAI-Bench. We prompt the existing MLLMs to evaluate the generated images and videos on GenAI-Bench and compute the accuracy with human voting. The experiment showed that the open-source MLLMs achieve very low performance, even the best model GPT-4o can only achieve $49.19\%$ accuracy. This is mostly because their lack of instruction-following and reasoning ability in complex vision scenarios. In the future, we will continue collecting human votes to update the leaderboard, helping the community to keep track of the research progress. We also plan to develop a more robust MLLM to better approximate human ratings in GenAI-Bench.